INTERNAL ROTATION IN THE NITROGEN COMPLEX OF 1,4-DIFLUOROBENZENE

\underline{MARTIN SCHÄFER}, ETH Zürich, Laboratorium für Physikalische Chemie, ETH-Hönggerberg HCI, CH-8093 Zürich, Switzerland.

The rotationally resolved S_{1} ← S_{0} spectrum of the 1,4-difluorobenzene–N_{2} van der Waals complex has been measured by Kang and Pratt^{a}. Due to internal motion, the spectrum is split into two subbands, of which only the stronger one can be fit with a rigid rotor Hamiltonian. The analysis of the stronger subband has revealed that N_{2} is located above the ring plane of 1,4-difluorobenzene. In order to obtain information about the internal motion, both subbands were analyzed with a semirigid C_{2v} frame–C_{2v} top internal rotor model^{b}. Upper limits of 20 cm^{−1} (S_{0}) and 6 cm^{−1} (S_{1}) for the barriers hindering N_{2} internal rotation have been obtained. Details about the problems in determing the equilibrium orientation and about implication on the structure determination will be discussed.

aCh. Kang and D. W. Pratt, 54th International Symposium on Molecular Spectroscopy, OSU, Columbus, paper WJ12, 1999.

bM. Schäfer, J. Chem. Phys. 115, 11139 (2001).